a)    Find the work done by the force field F on a particle that moves along the curve C. F(x, y) = (x2 + xy)i + (y – x2 y)j               C : x = t, y = 1/t (1 ≤ t ≤ 3)

b)    Use Green’s Theorem to evaluate the integral, assume that the curve C is oriented counterclockwise.  , where C is the triangle with vertices      (0, 0), (3, 3), and (0, 3).

c)     Evaluate the surface integral   , where σ is the part of the cone z =x2 + y2 that lies between the planes z = 1 and z = 2.

d)    Use the Divergence Theorem to find the flux of F across the surface σ with outward orientation.   F(x, y, z) = z3 i – x3 j + y3 k, where σ is the sphere x2 + y2 + z2 = a2.


1
Expert's answer
2021-02-01T05:46:42-0500

a) "W=\intop_CF\cdot dr"

"r(t)=(t,1\/t)"

"r'(t)=(1,-1\/t^2)"

"F(t)=(t^2+1,\frac{1}{t}-t)"

"\displaystyle\intop_1^3(t^2+1,\frac{1}{t}-t)\cdot(1,-1/t^2)dt=\displaystyle\intop_1^3(t^2+1-\frac{1}{t^3}+\frac{1}{t})dt="

"=(\frac{t^3}{3}+t+\frac{1}{2t^2}+lnt)|^3_1=9+1-\frac{1}{27}+\frac{1}{3}-1-1+1-1=8+\frac{8}{27}=\frac{224}{27}"


b)

"\oint_C(x^2+xy)dx+(y-x^2y)dy=\iint_D(Q_x-P_y)dA="

"=\iint_D(-2xy-x)dA=\displaystyle\intop_0^3dy\displaystyle\intop_0^y(-2xy-x)dx="

"=-\displaystyle\intop_0^3(y^3+\frac{y^2}{2})dy=-(\frac{y^4}{4}+\frac{y^3}{6})|^3_0=-(\frac{81}{4}+\frac{27}{6})=-\frac{297}{12}"


c) Parametri equations of the cone:

"x=zcost, y=zsint, z=z" ("0\leq t\leq 2\pi, 1\leq z \leq 2" )

"r(t,z)=zcosti+zsintj+zk"

"r_t\times r_z=\begin{vmatrix}
 i & j&k \\
 -zsint & zcost&0 \\ 
 cost&sint&1 \\
\end{vmatrix}=zcosti+zsintj-zk"

"|r_t\times r_z|=z\sqrt{2}"

"\iint_SdS=\iint_Q|r_t\times r_z|dA=\sqrt{2}\intop_0^{2\pi}dt\intop_1^2zdz=3\pi\sqrt{2}"


d)

"\iint_SF\cdot dS=\iiint_CdivFdV"

"divF=\frac{\partial F_x}{\partial x}+\frac{\partial F_y}{\partial y}+\frac{\partial F_z}{\partial z}"

"divF=0"

"\iint_SF\cdot dS=0"